OpenStudy (anonymous):

prob==> Solve by factoring: 4m^2+7m=0 Answer is ==> m(4m+7) but how do I solve to allow each factor to zero? please show me. thanks

Parth (parthkohli):

Now use the zero-product rule. \( \color{Black}{\Rightarrow m = 0 }\) \( \color{Black}{\Rightarrow 4m + 7 = 0 }\) Solve these two ^ and you get your two solutions I believe ^.^

OpenStudy (anonymous):

i don't get what you mean. you let the factors equal to zero so you get m = 0 and 4m + 7 = 0. solve for m for both cases.

OpenStudy (anonymous):

where are you getting confused?

OpenStudy (anonymous):

yes getting there confused

OpenStudy (anonymous):

you know how m(4m+7) was got right?

Parth (parthkohli):

Umm, so now is it easy for you to do it?

OpenStudy (anonymous):

if we have: a.b = 0; so we have: a = 0 or b = 0; so on this case: m(4m+7) = 0, you can let m = a and 4m+7 = b; so we have m=0 and 4m+7=0 so is it easy for you now?

OpenStudy (anonymous):

I understand now seeings how you made it a=0 etc. thanks it makes sence now

Parth (parthkohli):

Why didn't you ask us if you wanted to understand this?